OpenStudy (wampominater):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin2 2x - 2 sin2 x

OpenStudy (wampominater):

looks like this \[\cos4x+\cos2x=2-2\sin^22x-2\sin^2x\]

OpenStudy (lynfran):

using the double angle identity for cos cos2x=-2sin^2(x)+1 and cos4x= -2sin^2(2x)+1 so cos4x+cos2x \[(-2\sin ^{2}2x+1)+(-2\sin ^{2}x+1)\]

OpenStudy (lynfran):

remove the brackets and u will get the right hand side...

OpenStudy (anonymous):

@LynFran sup

OpenStudy (wampominater):

alright, Thank you!

OpenStudy (anonymous):

@LynFran sup its me Peachesjpg I changed my name y u blocked me